1.A bank operates with reserves of $100, loans of $300 and securities of $100. The bank’s only liability is deposits of $400 since it has zero debt. Calculate the bank’s leverage ratio.
Answer #1

ANSWER:

Assets = reserves + loans + securities = 100 + 300 + 100 = 500

debt = 400

leverage ratio = debt / assets = 400 / 500 = 0.8

so the bank's leverage ratio is 0.8